A pendulum clock gives correct time at 20 degree Celsius at a place where g= 9.800 m/s^2. The pendulum consists of a light steel rod connected to a heavy ball. It is taken to a different place where g=9.788 m/s^2. At what temperature will it give correct time?coefficient of linear expansion of steel = 12x 10^-6/ degree celsius.


Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Cathy,

Ans: The time period is T=2∏√(L/g)

So if the time period is the same then we get,

T=constant hence L1/L2=g1/g2

Now as the value of g is decreasing and hence temp must be decreased to get the same value of T. Now L2=L1(1-αt)

solving we get,

t= -122 degree

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Cathy !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur